# LittleBit: Ultra Low-Bit Quantization via Latent Factorization

### v4

@@ A.2.1 Residual Compensation
As shown in  Fig.˜7, residual compensation preserves a high‑fidelity resemblance to the full‑precision baseline across all inspected layers. In the no-residual setting, the output feature distributions exhibit noticeable deformation relative to the full-precision baseline. In contrast, applying residual compensation consistently restores structural similarity to the original distribution, preserving both directional patterns and magnitude diversity.
@@
A.2.2 SmoothSign versus Straight-Through Estimator
@@
A.2.3 Grouped Query Attention
@@ A.3 Perplexity on C4 and PTB
A.4 Average Bits Per Weight Calculation
@@ A.5 Pruning versus SVD
A.6 Analysis of Generated Samples
@@
    Impact on Specificity, Factual Detail, and Coherence: As the effective bits decreases from 0.8 BPW towards 0.1 BPW, the generated text exhibits significant changes in detail, factual accuracy, and coherence. At 0.8 BPW, the Mona Lisa sample, while attempting to provide details (e.g., "complex facial features"), becomes verbose and repetitive, and anachronistically refers to Da Vinci and the Mona Lisa in the context of the "20th century." The Turing sample at this effective bit, however, offers a standard, factually sound definition. As the effective bit drops to 0.55 BPW, the Mona Lisa sample loses coherence significantly, expressing uncertainty about the painting’s location and bizarrely shifting to a help-seeking persona ("Could you help me find the Mona Lisa painting…"). The Turing sample at 0.55 BPW begins to broaden its definition, losing some precision but remaining somewhat related.

### v5

@@ B.1 Residual Compensation
As shown in  Fig.˜7, residual compensation preserves a high-fidelity resemblance to the full-precision baseline across all inspected layers. In the no-residual setting, the output feature distributions exhibit noticeable deformation relative to the full-precision baseline. In contrast, applying residual compensation consistently restores structural similarity to the original distribution, preserving both directional patterns and magnitude diversity.
@@
B.2 SmoothSign versus Straight-Through Estimator
@@
B.3 Grouped Query Attention
@@ Appendix C Perplexity on C4 and PTB
Appendix D Average Bits Per Weight Calculation
@@ Appendix E Pruning versus SVD
Appendix F Analysis of Generated Samples
@@
    Impact on Specificity, Factual Detail, and Coherence: As the effective bits decrease from 0.8 BPW towards 0.1 BPW, the generated text exhibits significant changes in detail, factual accuracy, and coherence. At 0.8 BPW, the Mona Lisa sample, while attempting to provide details (e.g., "complex facial features"), becomes verbose and repetitive, and anachronistically refers to Da Vinci and the Mona Lisa in the context of the "20th century." The Turing sample at this effective bit, however, offers a standard, factually sound definition. As the effective bit drops to 0.55 BPW, the Mona Lisa sample loses coherence significantly, expressing uncertainty about the painting’s location and bizarrely shifting to a help-seeking persona ("Could you help me find the Mona Lisa painting…"). The Turing sample at 0.55 BPW begins to broaden its definition, losing some precision but remaining somewhat related.